I'm trying to solve by elimination but the answers just don't seem right?
#1) 5x-7y=-16 10x-14y=32 5x-7(-7.90)=-16
2x-5y=11 10x+25y=-55 5x+55.3=-16
11y=-87 -55.3 -55.3 (-14.26,-7.90)
y=-7.90 5/5x = -71.3/5
x = -14.26
#2) 5x-9=7 7y-3x=-5
+9 +9 7y-3(16/5)=-5
5/5x=16/5 7y-42=-5 (16/5, 37/7)
x = 16/5 +42 +42
7/7y = 37/7
y = 37/7
#3) 9x+5y=6 2(3/11)-5y=-3
2x-5y=-3 6/11-5y=-3
11/11x = 3/11 -6/11 -6/11
x = 3/11 -5/-5y=-27/-5
y = -27/-5
#4) 8x-3y=16 (8/16,-4)
8x+3y=-8
Thanks.

Answer by Alan3354(69443)

5x-7y=-16 Eqn 1
10x-14y=-32 Eqn 1 times 2
2x-5y=11 Eqn 2
10x-25y=55 Eqn 2 times 5
10x-14y=-32 Eqn 1 times 2
------------------------- Subtract
-11y = 87
y = -87/11
-------------
2x + 435/11 = 11 Eqn 2
2x = -314/11
x = -157/11
------------
------------
11y=-87 -55.3 -55.3 (-14.26,-7.90)
y=-7.90 5/5x = -71.3/5
x = -14.26
--------------
#2) 5x-9=7 7y-3x=-5
+9 +9 7y-3(16/5)=-5
5/5x=16/5 7y-42=-5 (16/5, 37/7)
x = 16/5 +42 +42 **************** 3*16 = 48, not 42
7/7y = 37/7
y = 37/7
#3) 9x+5y=6 2(3/11)-5y=-3
2x-5y=-3 6/11-5y=-3
11/11x = 3/11 -6/11 -6/11
x = 3/11 -5/-5y=-27/-5 ******* -3 - 6/11 = -39/11
y = -27/-5 ****************** y = 39/55
-------------
#4) 8x-3y=16 (8/16,-4) --> (1/2,-4)
8x+3y=-8